Consider the following five bonds, all with notional amounts of $100.00, that are trading in a...

  1. Consider the following five bonds, all with notional amounts of $100.00, that are trading in a liquid market on September 30th 2018
    1. T-bill 1: 1 year maturity, no annual coupon, market price = $99.01
    2. T-bill 1: 3 year maturity, no annual coupon, market price = $92.86
    3. Bond 1: 4 year maturity, 4% annual coupon, market price = $103.92
    4. Bond 2: 5 year maturity, 2% annual coupon, market price = $95.52
    5. Bond 3: 5 year maturity, 8% annual coupon, market price = $123.31

Assume that all coupon and maturity dates align with all cashflows (coupons and notional amounts) paying at September 30th of each year.

Given this market, what are the discount factors and implied zero coupon rates associated with annual terms 1 – 5 years? (show all your work) …10 marks
Answer #1

Answer:

zero coupon rates:

i. T-bill 1: 1 year maturity, no annual coupon, market price = $99.01

We will use excel function RATE to get zero coupon rates:

RATE (nper, pmt, pv, fv, 0)

=RATE (1, 0, 99.01, -100, 0)

=1.00%

Zero coupon rates = 1.00%

ii. T-bill 1: 3 year maturity, no annual coupon, market price = $92.86

= RATE (3, 0, 92.86, -100, 0)

=2.50%

Zero coupon rates = 2.50%

Bonds:

iii. Bond 1: 4 year maturity, 4% annual coupon, market price = $103.92

Annual coupon = 100 * 4% = 4

Discount factor = RATE (4, 4, -103.92, 100, 0)

=2.95%

Discount factor = 2.95%

iv. Bond 2: 5 year maturity, 2% annual coupon, market price = $95.52

Annual coupon = 100 * 2% = 2

Discount factor = RATE (5, 2, -95.52, 100, 0) = 2.98%

Discount factor = 2.98%

v. Bond 3: 5 year maturity, 8% annual coupon, market price = $123.31

Annual coupon = 100 * 8% = 8

Discount factor = RATE (5, 8, -123.31, 100, 0) = 2.92%

Discount factor = 2.92%